# A Gaussian smooth transition vector autoregressive model: An application to the macroeconomic effects of severe weather shocks

### v2

@@ 2.3 Gaussian STVAR model
2.3.1 Comparison to related models
Our GSTVAR model has a number of desirable features compared to popular alternative nonlinear VAR models. As already discussed, in line with the GMVAR model of Kalliovirta et al. (2016), the transition between the regimes depends on multiple past observations in contrast to the Markov-switching VAR (MS-VAR) model, where the transition probabilities depend on the most recent regime, or the logistic STVAR (LSTVAR) and threshold VAR (TVAR) models, where the regime is determined by the lagged value of an observable variable or an exogenous switching variable. On the other hand, like the LSTVAR model, the GSTVAR model facilitates gradual shifts between the regimes, whereas in the TVAR, GMVAR and MS-VAR models each observation is generated from a single regime. Hence, the GSTVAR model combines the advantages of the GMVAR, TVAR, LSTVAR and MS-VAR models.
Although the transition weight function of the GSTVAR model is similar to the mixing weights of the GMVAR model, it differs from that model in other ways. Specifically, the GMVAR model generates each observation from a randomly selected distinct regime, while the GSTVAR model allows for smooth transitions between the regimes and thereby facilitates gradual shifts in the dynamics. This is the case because in the GMVAR model, the transition weights $\alpha_{m,t}$ are replaced by unobservable regime variables $s_{m,t}$ in Equations (2.1)–(2.3). These regime variables $s_{m,t}$, $m=1,...,M$, are such that at each $t$, exactly one of them takes the value one, whereas the rest take value zero according to the probabilities $P(s_{m,t}=1|\mathcal{F}_{t-1})=\alpha_{m,t}$. Because the regime that generates each observation is thus unknown in the GMVAR model, the (structural) errors cannot, in general, be recovered from the data, while they are readily available in the GSTVAR model. This is particularly important in structural analysis because it enables computing the historical decomposition and conducting counterfactual analyses.
$\alpha_{1,t}=1-\alpha_{2,t},\ \ \alpha_{2,t}=[1+\exp\{-\gamma(y_{i,t-j}-c)\}]^{-1},$ (2.10)
where $y_{i,t-j}$ is the $j$th lagged observation $(j\in\{1,....,p\})$ of the $i$th variable $(i\in\{1,....,d\})$, $c\in\mathbb{R}$ is a location parameter, and $\gamma>0$ is a scale parameter. The location parameter c determines the mid-point of the transition function, i.e., the value of the switching variable when the weights are equal. The scale parameter $\gamma$, in turn, determines the smoothness of the transitions (smaller $\gamma$ implies smoother transitions), and it is assumed strictly positive so that $\alpha_{2,t}$ is increasing in $y_{it-j}$. This simple setup can be generalized in various directions (see, e.g., Teräsvirta and Yang, 2014). It is also possible to use an exogenous or deterministic switching variable instead of a lagged endogenous variable in the logistic transition weights (2.10). However, in these cases, it may be difficult to determine whether the model is stationary, and it is not generally possible to estimate the true nonlinear generalized impulse response functions.
The logistic transition weights have the advantage that the regimes typically have clear interpretations and the switching variable can be defined to address some specific aspect of the empirical problem. For example, when using the lagged GDP growth as the switching variable, the regimes can be interpreted as recession and expansion. In comparison, the transition weights of the GSTVAR model depend on the full distribution of the preceding $p$ observations and can thereby capture more complicated switching-dynamics, which facilitates associating the different regimes with different aspects of the phenomenon being modelled. As the transition weights of the GSTVAR model are not affected by the choice of a switching variable but determined based on the weighted relative likelihoods of the regimes, they are more clearly data-driven and their own evolution is interesting as such. However, the resulting regimes are statistically determined, so it is not always clear how to interpret them in terms of a given application.
$\alpha_{m,t}=I(c_{m-1}<s_{t}<c_{m}),\ \ m=1,...,M,$ (2.11)
$P(s_{m,t}=1|s_{n,t-1}=1)=p_{mn}\ \text{for all}\ t\ \text{and}\ m,n=1,...,M,$ (2.12)
that satisfy $\sum_{n=1}^{M}p_{mn}=1$ for all $m$. In other words, the regime that generates each observation is selected randomly at each $t$ according to the probabilities $p_{mn}$, $m,n=1,...,M$. While the MS-VAR model is a relatively simple framework for flexibly modelling nonlinear time series with unobservable discrete regime-switches, unlike our GSTVAR model, it is unable to capture gradual shifts in the dynamics of the data. Moreover, in the same way as in the GMVAR model, it is not possible to recover structural shocks from the fitted model because the regime that generates each observation is not observed.
@@ 3.1 Maximum likelihood estimation
The parameters of the GSTVAR model can be estimated by the method of maximum likelihood (ML). We collect the parameters to the vector
$\boldsymbol{\theta}=(\phi_{1,0},...,\phi_{m,0},\varphi_{1},...,\varphi_{M},\sigma,\alpha),$ (3.1)
where $\varphi_{m}=(\text{vec}(A_{m,1}),....,\text{vec}(A_{m,p}))$, $m=1,...,M$, $\sigma=(\text{vech}(\Omega_{1}),...,\text{vech}(\Omega_{M}))$, $\alpha=(\alpha_{1},...,\alpha_{M-1})$ contains the transition weight parameters (notice that $\alpha_{M}=1-\sum_{m=1}^{M-1}\alpha_{m}$).
@@
$L_{t}(\boldsymbol{\theta})=\sum_{t=1}^{T}l_{t}(\boldsymbol{\theta})=\sum_{t=1}^{T}\log n_{d}(y_{t};\mu_{y,t},\Omega_{y,t}).$ (3.2)
@@
To ensure ergodic stationarity of the process, we assume that Condition 1 in Section 2.1 holds. Moreover, it is assumed that the true parameter value is an interior point of a compact subset of the parameter space, which is a standard condition for asymptotic normality of the ML estimator. Thus, given ergodic stationarity of the process, there is no particular reason to believe that the standard asymptotic results of consistency and asymptotic normality would not apply to the ML estimator. Finally, to achieve identification of the parameters such that the regimes cannot ’relabelled’ to obtain the same model with different parameter vectors,we order the transition weight parameters $\alpha_{m}$, $m=1,...,M$ in a decreasing order. That is, we assume
@@
The true parameter value $\boldsymbol{\theta}_{0}$ is an interior point of $\boldsymbol{\Theta}$, which is a compact subset of the parameter space $\{\boldsymbol{\theta}=(\phi_{1,0},...,\phi_{m,0},\varphi_{1},...,\varphi_{M},\sigma,\alpha,\nu)\in\mathbb{R}^{M(d+d^{2}p+d(d+1)/2)}\times(0,1)^{M-1}\times(2,\infty):$ $\Omega_{m}$ is positive definite for all $m=1,...,M$, and Conditions 1 and (3.3) hold.}}\}}.
Finding the ML estimate amounts to maximizing the log-likelihood function in (3.2) over a high dimensional parameter space satisfying the constraints in Assumption 1. Due to the complexity of the log-likelihood function, numerical optimization methods are required. The maximization problem can, however, be challenging in practice due to the complicated dependence of the transition weights on the preceding observations, which induces a large number of modes to the surface of the log-likelihood function, and large areas to the parameter space, where it is flat in multiple directions.
We follow Virolainen (2022, forthcoming) and Meitz et al. (2023) and employ a two-phase estimation procedure that is ran for a large number of times. In the first phase, a genetic algorithm is used to find parameter values that are near local maximum points of the log-likelihood function. Since genetic algorithms tend to converge slowly near local solutions, a gradient based variable algorithm is ran for each of the starting values, resulting in a number of alternative local solutions. Some of the estimation rounds may end up in saddle points or near-the-boundary points that are not local solutions, and some of the local solutions may be inappropriate for statistical inference (for instance, there might be only few observations from some of the regimes). Because Condition 1 included in Assumption 1 is computationally costly to verify, we recommend using the necessary Condition 2 to restrict the parameter space.
After the estimation rounds have been ran, the researcher can choose among the appropriate local solutions the one that maximizes the log-likelihood. Then, the sufficient Condition 1 for ergodic stationarity can be checked for the selected local solution if the more easily verified necessary Condition 2 is used to restrict the parameter space instead of Condition 1 in the estimation. The accompanying R package sstvars (Virolainen, 2024) employs a modified genetic algorithm that works similarly to the one described in Virolainen (2022).
@@ 3.2 Model selection
It is well known that testing for linearity against a model with multiple regimes, in general, poses a nonstandard testing problem because the model is identified only under the alternative (Davies, 1977). This is the case also in our setup, and, therefore, we follow the previous literature and recommend using information criteria and residual-based diagnostic checks to compare the fit of models with different orders and numbers of regimes, as well as to study their adequacy in capturing the autocorrelation structure, conditional heteroskedasticity, and distribution of the data. The derivation of formal diagnostic tests is beyond the scope of this paper, but graphical devices, including quantile-quantile plots as well as autocorrelation and cross-correlation functions of the residuals and their squares, can be used to this end.
For a number of reasons, GSTVAR models of a relatively low autoregressive order are preferable. Firstly, the estimation of GSTVAR models with a large $p$ can be tedious in practice. Secondly, if the order $p$ is large, the number of parameters increases vastly when the number of regimes is increased, which substantially decreases estimation accuracy. Thirdly, unlike in linear VARs, decreasing the autoregressive order may actually improve the fit because the transition weights are calculated using the whole joint distribution of the preceding $p$ observations. Consequently, the sensitivity of the transition weights to individual observations and, hence, responsiveness to changes in the dynamics of the data decrease with $p$.
It is also advisable to be conservative with the number of regimes $M$ because if the number of regimes is chosen too large, some of the parameters in the model are not identified (see the related discussion in Kalliovirta et al., 2016, Section 3.4). Moreover, increasing the number of regimes substantially increases the complexity of the surface of the log-likelihood function, making the estimation of the parameters challenging in practice.
@@ 4 Empirical application
In this section, we apply our GSTVAR model to study the effects of severe weather on the U.S. macroeconomy. This issue has previously been addressed by Kim et al. (2022), who employ a two-regime recursive STVAR model with transition weights defined by a linear time trend ($\alpha_{2,t}=t/T$ and $\alpha_{1,t}=1-\alpha_{2,t}$). They argue that allowing for time-varying coefficients in the model is important because the economic effects of severe weather may have changed due to actions potentially taken for the economy to adapt to climate change. According to Kim et al. (2022), there is little evidence in favor of adaptation to severe weather in the U.S, as the impacts of the severe weather shocks are found significant only towards the end of the sample (spanning from 1963 to 2019).
The linear time trend might not, however, sufficiently describe the adaptation of the economy, because the effects of severe weather may vary across different states of the economy due to variation in factors such as fiscal flexibility and sector-specific vulnerabilities. It is not obvious how such variation should be accommodated in the transition weights in a nonlinear SVAR model. We address this issue with our GSTVAR model in which the regimes and shifts between them are formed based on the statistical properties of the data through the full distribution of the preceding $p$ observations. This approach plausibly better describes the evolution of the joint dynamics of severe weather and U.S. macroeconomy through time than an STVAR model with transition weights defined by a linear time trend.
@@ 4.1 Data
We consider a four-variable monthly U.S. data set that covers the period from 1961:1 to 2022:3 ($T=735$ observations) and includes indicators for severe weather and economic activity as well as consumer price inflation and an interest rate variable. As argued by Kim et al. (2022), it is important to use data of more high frequency than quarterly because some weather effects can be short-lived, and the end point of our sample period is determined by the availability of the monthly GDP growth data. While the sample period is slightly longer than in Kim et al. (2022), our general conclusions are not affected by using their shorter sample period (from 1963:4 to 2019:5), as discussed in Appendix C.2.
As an indicator of the frequency of severe weather and the extent of sea level rise, we use the Actuaries Climate Index (ACI), developed by actuarial organizations in the United States and Canada (American Academy of Actuaries et al., 2023). Following Kim et al. (2022), we seasonally adjust the ACI series with the standard Census Bureau X-13 seasonal adjustment algorithm. For a more detailed description of ACI, see Kim et al. (2022).
As an aggregate measure of real economic activity, we use the monthly GDP growth rate constructed by the Federal Reserve Bank of Chicago from a collapsed dynamic factor analysis of a panel of $500$ monthly measures of real economic activity and quarterly real GDP growth (Federal Reserve Bank of Chicago, 2023). While Kim et al. (2022) use both the industrial production index and unemployment rate to measure real economic activity at the monthly frequency, we opt for the monthly GDP growth rate instead. It has the advantage of being a more comprehensive measure of real economic activity, including agricultural output, tourism, and services, which can be significantly affected by the severe weather shocks.
Finally, we include the monthly growth rate of the consumer price index (CPI) and the effective federal funds rate that is replaced by the Wu and Xia (2016) shadow rate for the zero lower bound periods. The CPI and federal funds rate series were downloaded from the Federal Reserve Bank of S.t. Louis database and Wu and Xia (2016) shadow rate from the Federal Reserve Bank of Atlanta database. The time series are presented in the top four panels of Figure 1, where the shaded areas indicate the U.S. recessions defined by the NBER.
@@ 4.2 GSTVAR model
To select the order of our GSTVAR model, we start by examining the partial autocorrelation functions of the U.S. time series (shown in Figure C.1 of Appendix C). They suggest that the autoregressive order $p=4$ could be reasonable for a parsimonious model. Therefore, we fit two-regime GSTVAR models with autoregressive orders $p=1,....,5$ and, to compare them, compute the values of three information criteria (AIC, BIC and HQIC). As seen in Table 1, the order $p=4$ minimizes the AIC, whereas the order $p=2$ minimizes the HQIC and BIC. The table also contains similar results for linear VAR (i.e., one-regime GSTVAR) models. In terms of the information criteria, they are clearly inferior to our two-regime GSTVAR models. While adding a third regime to the model might improve the fit, the number of parameters would be large compared the number of observations from each regime, substantially decreasing estimation accuracy. Moreover, since incorporating too many regimes in the model would also result in identification issues (see Section 3.2), we confine ourselves to two-regime models.
@@
Table 1: The values of the log-likelihood function and the information criteria divided by the number of observations ($735-p$) for a number of linear VAR and two-regime GSTVAR models.
@@
Based on graphical residual diagnostics, presented in Appendix C, the two-regime GSTVAR model with $p=4$ lags captures the autocorrelation structure of the data reasonably well. Some of the conditional heteroskedasticity and marginal distribution of the series are not captured, but the inadequacies are not particularly severe. Finally, the sufficient Condition 1 for ergodic stationarity holds for the selected model, which is seen by finding an upper bound of the joint spectral radius of the matrices $\boldsymbol{A}_{m}$ (2.4), $m=1,2$. with the JSR toolbox (Jungers, 2023) in MATLAB. The resulting upper bound is $0.995$, which is strictly smaller than one, satisfying the stationary condition. Hence, we proceed with the two-regime fourth-order model.
@@
The estimated transition weights indicate the relative importance of each regime, and the time series of the weights of Regime 1 are presented in the bottom panel of Figure 1. The corresponding weights of Regime 2 are, of course, obtained by subtracting these weights from unity. It is not straightforward to give the regimes an economic interpretation. Regime 1 mainly dominates in the 1960’s, 1970’s, and 1980’s, but obtains large weights also during the Financial crisis and from the beginning of the COVID-19 crisis onwards (excluding a short period in 2021). Overall, Regime 1 thereby mostly prevails in the earlier sample and Regime 2 in the later sample. The dominance of Regime 1 during the volatile periods of 1970’s and 1980’s as well as during the Financial crisis and the COVID-19 crisis, suggests that it could, to some extent, represent more turbulent times compared to the second regime.

### v3

@@ 2.3 Gaussian STVAR model
Our GSTVAR model has a number of desirable features compared to popular alternative nonlinear VAR models. As already discussed, in line with the GMVAR model of Kalliovirta et al. (2016), the transition between the regimes depends on the full distribution of multiple past observations. By contrast, in the Markov-switching VAR (MS-VAR) model,the transition probabilities depend only on the most recent regime, whereas in the logistic STVAR (LSTVAR) and threshold VAR (TVAR) models, the regime is determined by the lagged value(s) of some observable variable(s) or exogenous switching variable(s). On the other hand, like the LSTVAR model, the GSTVAR model facilitates gradual shifts between the regimes, whereas in the TVAR, GMVAR and MS-VAR models each observation is generated from a single regime. Hence, the GSTVAR model combines several advantages of the GMVAR, TVAR, LSTVAR and MS-VAR models. A more detailed comparison of the models is given in Appendix C.
@@ 3.1 Maximum likelihood estimation
The parameters of the GSTVAR model can be estimated by the method of maximum likelihood (ML). We collect the parameters to the vector $\boldsymbol{\theta}=(\phi_{1,0},...,\phi_{m,0},\varphi_{1},...,\varphi_{M},\sigma,\alpha)$, where $\varphi_{m}=(\text{vec}(A_{m,1}),....,\text{vec}(A_{m,p}))$, $m=1,...,M$, $\sigma=(\text{vech}(\Omega_{1}),...,\text{vech}(\Omega_{M}))$, and $\alpha=(\alpha_{1},...,\alpha_{M-1})$ contains the transition weight parameters (notice that $\alpha_{M}=1-\sum_{m=1}^{M-1}\alpha_{m}$).
@@
$L_{t}(\boldsymbol{\theta})=\sum_{t=1}^{T}\log n_{d}(y_{t};\mu_{y,t},\Omega_{y,t}).$ (3.1)
@@
To ensure ergodic stationarity of the process, we assume that Condition 1 in Section 2.1 holds. Moreover, it is assumed that the true parameter value is an interior point of a compact subset of the parameter space, which is a standard condition for asymptotic normality of the ML estimator. Thus, given ergodic stationarity of the process, there is no particular reason to expect that the standard asymptotic results of consistency and asymptotic normality would not apply to the ML estimator. Finally, to achieve identification of the parameters such that the regimes cannot ’relabelled’ to obtain the same model with different parameter vectors,we order the transition weight parameters $\alpha_{m}$, $m=1,...,M$ in a decreasing order. That is, we assume
@@
The true parameter value $\boldsymbol{\theta}_{0}$ is an interior point of $\boldsymbol{\Theta}$, which is a compact subset of the parameter space $\{\boldsymbol{\theta}=(\phi_{1,0},...,\phi_{m,0},\varphi_{1},...,\varphi_{M},\sigma,\alpha)\in\mathbb{R}^{M(d+d^{2}p+d(d+1)/2)}\times(0,1)^{M-1}:$ $\Omega_{m}$ is positive definite for all $m=1,...,M$, and Conditions 1 and (3.2) hold.}}\}}.
Finding the ML estimate amounts to maximizing the log-likelihood function in (3.1) over a high dimensional parameter space satisfying the constraints in Assumption 1. Due to the complexity of the log-likelihood function, numerical optimization methods are required. The maximization problem can, however, be challenging in practice due to the complicated dependence of the transition weights on the preceding observations, which induces a large number of modes of the log-likelihood function and large areas of the parameter space, where it is flat in multiple directions.
We follow Virolainen (2022, 2025b) (and others) and employ a two-phase estimation procedure that is run for a large number of times. [^2] For example, in our empirical application presented in Section 4, we use $6000$ estimation rounds to estimate the four-variate GSTVAR with $p=4$, $M=2$. In the first phase, a genetic algorithm is used to find parameter values that lie close to the local maximum points of the log-likelihood function. Since genetic algorithms tend to converge slowly near local solutions, a gradient based variable metric algorithm is run for each of the starting values, resulting in a number of alternative local solutions. Some of the estimation rounds may end up in saddle points or near-the-boundary points that are not local solutions, and some of the local solutions may be inappropriate for statistical inference (for instance, there might be only few observations from some of the regimes). Because Condition 1 included in Assumption 1 is computationally costly to verify, we recommend using the necessary Condition 2 to restrict the parameter space.
After the estimation rounds have been run, the researcher can choose among the appropriate local solutions the one that maximizes the log-likelihood. Then, it can be verified that the selected local solution satisfies the sufficient Condition 1 for ergodic stationarity if the more easily checked necessary Condition 2 is used to restrict the parameter space in estimation. The accompanying R package sstvars (Virolainen, 2025a) employs a modified genetic algorithm that works similarly to that described in Virolainen (2022).
To study the finite-sample properties of the ML estimator, we conduct a small scale Monte Carlo study, which is discussed in detail in Appendix D. Because estimation is computationally demanding, we consider a simple bivariate GSTVAR model with $M=2$ and $p=1$. We consider two specifications that differ with respect to the values of the autoregressive parameters and intercepts (see Table D.1 in Appendix D). According to the results (see Table D.2 in Appendix D) the estimator is slightly biased in small samples, but the bias vanishes and estimation accuracy increases with the sample size. Also, estimation accuracy does not seem to depend on the specification.
@@ 3.2 Model selection
It is well known that testing for linearity against a model with multiple regimes, in general, poses a nonstandard testing problem because the model is identified only under the alternative (Davies, 1977). This is the case also in our setup, and, therefore, we follow the previous literature (e.g., Kalliovirta et al., 2016 and Virolainen, 2022, 2025b) and recommend using information criteria and residual-based diagnostic checks to compare the fit of models with different orders and numbers of regimes, as well as to study their adequacy in capturing the autocorrelation structure, conditional heteroskedasticity, and distribution of the data. The derivation of formal diagnostic tests is beyond the scope of this paper, but graphical devices, including quantile-quantile plots as well as autocorrelation and cross-correlation functions of the residuals and their squares, can be used to this end.
For a number of reasons, GSTVAR models of a relatively low autoregressive order are preferable. Firstly, the estimation of GSTVAR models with a large $p$ can be tedious in practice. Secondly, if the order $p$ is large, the number of parameters increases vastly when the number of regimes is increased, which is likely to substantially decrease estimation accuracy. Thirdly, unlike linear VARs, decreasing the autoregressive order may actually improve the fit because the transition weights are calculated using the whole joint distribution of the preceding $p$ observations. Consequently, the sensitivity of the transition weights to individual observations and, hence, responsiveness to changes in the dynamics of the data decrease with $p$.
It is also advisable to be conservative with the number of regimes $M$ because if the number of regimes is chosen too large, some of the parameters in the model are not identified (see the related discussion in Kalliovirta et al., 2016, Section 3.4). Having too many regimes may also lead to overfitting, as the number of parameters may become overly large compared to the number of observations in some of the regimes. Moreover, increasing the number of regimes substantially increases the complexity of the surface of the log-likelihood function, making estimation of the parameters challenging in practice.
@@ 4 Empirical application
In this section, we apply our GSTVAR model to study the effects of severe weather on the U.S. macroeconomy. This issue was recently addressed by Kim et al. (in press), who employ a two-regime recursive STVAR model with transition weights defined by a linear time trend ($\alpha_{2,t}=t/T$ and $\alpha_{1,t}=1-\alpha_{2,t}$ in (2.1)). They argue that allowing for time-varying coefficients in the model is important because the strength of the economic impact of severe weather may have diminished due to actions potentially taken to make the economy more adaptable to climate change. However, they find little evidence in favor of adaptation to severe weather in the U.S, but on the contrary, their results suggest that the severe weather shock has a significant effect only towards the end of the sample (spanning from 1963 to 2019).
Using transition weights defined by a linear time trend in a two-regime STVAR model only allows one smooth transition from one regime to the other during the sample period. This approach may be too simple to describe the adaptation of the economy, as the effects of severe weather may vary across different states of the economy due to variation in factors such as fiscal flexibility and sector-specific vulnerabilities. However, it is not obvious which exogenous or lagged endogenous variables would be capable of capturing such variation in the transition weights in a nonlinear SVAR model. Therefore, we address this issue with our GSTVAR model in which the regimes and shifts between them are formed based on the statistical properties of the data through the full distribution of the preceding $p$ observations. Due its more data-driven nature, this approach plausibly better describes the evolution of the joint dynamics of severe weather and U.S. macroeconomy through time than an STVAR model with transition weights defined by a linear time trend or depending only on the level of exogenous or lagged endogenous variables.
@@ 4.1 Data
We consider a four-variable monthly U.S. data set that comprises indicators of severe weather and economic activity as well as consumer price inflation and an interest rate variable. While our sample period (from 1961:1 to 2022:3, $T=735$ observations) is somewhat longer than than in Kim et al. (in press), the main conclusions based on their shorter sample period (from 1963:4 to 2019:5) remain the same (see Appendix E.2.1 for the subsample results). The end point of our sample period is determined by the availability of the monthly GDP growth data, but as argued by Kim et al. (in press), it is important to use data of higher frequency than quarterly because some weather effects can be short-lived.
As an indicator of the frequency of severe weather and the extent of sea level rise, we use the Actuaries Climate Index (ACI), developed by actuarial organizations in the United States and Canada (American Academy of Actuaries et al., 2023). Following Kim et al. (in press), we seasonally adjust the ACI series with the standard Census Bureau X-13 seasonal adjustment algorithm. For a more detailed description of ACI, see Kim et al. (in press).
As an aggregate measure of real economic activity, we use the monthly GDP growth rate constructed by the Federal Reserve Bank of Chicago from a collapsed dynamic factor analysis of a panel of $500$ monthly measures of real economic activity and quarterly real GDP growth (Federal Reserve Bank of Chicago, 2023). While Kim et al. (in press) use both the industrial production index and unemployment rate to measure real economic activity at the monthly frequency, we opt for the monthly GDP growth rate instead. It has the advantage of being a more comprehensive measure of real economic activity, including agricultural output, tourism, and services, which can be significantly affected by the severe weather shocks. Moreover, using only a single variable to measure real economic activity is likely to facilitate estimation by reducing the dimension, and thereby also the number of parameters, of the model.
Finally, we include the monthly growth rate of the consumer price index (CPI) and the effective federal funds rate (RATE) that is replaced by the Wu and Xia (2016) shadow rate for the zero lower bound periods. The CPI and federal funds rate series were downloaded from the Federal Reserve Bank of S.t. Louis database and Wu and Xia (2016) shadow rate from the Federal Reserve Bank of Atlanta database. The time series are depicted in the top four panels of Figure 1, where the shaded areas indicate the U.S. recessions defined by the NBER.
@@ 4.2 GSTVAR model
To select the order of the GSTVAR model, we start by examining the partial autocorrelation functions of the U.S. time series (shown in Figure E.1 of Appendix E). They suggest that the autoregressive order $p=4$ might be adequate. Therefore, we fit two-regime GSTVAR models with autoregressive orders $p=1,....,5$ and, to compare them, compute the values of three information criteria (AIC, BIC and HQIC). As seen in Table 1, the order $p=4$ minimizes the AIC, whereas the order $p=2$ minimizes the HQIC and BIC. The table also contains similar results for linear VAR (i.e., GSTVAR with $M=1$) models, which are clearly inferior to the two-regime GSTVAR models in terms of the information criteria. Moreover, the constancy of the AR matrices and intercepts as well as the constancy of AR matrices only are clearly rejected by the Wald test (see Appendix E). While adding a third regime to the model might improve the fit, the number of parameters compared the number of observations in each regime would increase substantially, possibly leading to overfitting. Moreover, since incorporating too many regimes in the model could also result in identification issues (see Section 3.2), we confine ourselves to two-regime models.
@@
Table 1: The values of the log-likelihood function and the information criteria (divided by the number of observations) for a number of linear VAR and two-regime GSTVAR models.
@@
Based on graphical residual diagnostics, presented in Appendix E, the two-regime GSTVAR model with $p=4$ lags captures the autocorrelation structure of the data reasonably well. Conditional heteroskedasticity and marginal distribution of the series are not completely captured, but the inadequacies are not particularly severe. Finally, the sufficient Condition 1 for ergodic stationarity holds for the selected model, as the upper bound of the joint spectral radius of the matrices $\boldsymbol{A}_{m}$ (2.4), $m=1,2$ (see Section 2.1), is found to be strictly less than one ($0.995$). Hence, we proceed with the two-regime fourth-order model.
@@
The estimated transition weights indicate the relative importance of each regime. The time series of the weights of Regime 1 are presented in the bottom panel of Figure 1, and the corresponding weights of Regime 2 are, of course, obtained by subtracting these weights from unity. The shifts between the regimes are relatively fast, with the switch from one regime to the other typically taking from one to three months to be completed. Regime 1 mainly dominates in the 1960s, 1970s, and 1980s, but obtains large weights also during the Financial crisis and from the beginning of the COVID-19 crisis onward (excluding a short period in 2021). Overall, Regime 1 thereby mostly prevails in the earlier sample and Regime 2 in the later sample. The stationary standard deviations of ACI, GDP growth rate, CPI growth rate, and interest rate in Regime 1 are $0.59,1.17,0.47$, and $9.57$, respectively, whereas they are $0.52,0.29,0.26$, and $3.46$ in Regime 2. These differences and the dominance of Regime 1 in the volatile periods of 1970s and 1980s as well as during the Financial crisis and the COVID-19 crisis, suggest that it could represent more turbulent times compared to Regime 2.
4.3 Structural analysis